### v1

@@ 4.2.1 Data and Resources
4.2.2 Procedure
Each agent in CT-Agent was tasked with specific roles, as outlined in the Methods section. The Safety Agent queried drug databases, the Efficacy Agent analyzed disease information, the Enrollment Agent predicted the enrollment difficulty, and the Planning Agent synthesized these findings into actionable insights. We conducted experiments to assess the accuracy of outcome predictions and the precision of failure reason identifications.
4.3 Implementation Details
@@
Each agent within the CT-Agent framework is designated a specific role, which is integrated directly into the LLM’s system prompt for clarity and focus. For instance, the role of the Efficacy Agent is defined as follows:
@@
These implementation strategies collectively ensure that each component of CT-Agent operates effectively and that the integration between different agents and external tools is seamless, fostering an environment conducive to robust, reproducible research.
4.4 Quantitative Results
@@
Table 2 presents the performance of various methods. Specifically, CT-Agent obtain the highest ROC-AUC score at 0.8347 among all the compared methods. We observe that CT-Agent achieve competitive performance among all the well-established methods, e.g., GBDT. Also, compared with the standard prompt (basic GPT) model, our method consistently improves all six evaluation metrics. We also compare the performance of GPT-3.5 and GPT-4 for the standard prompting method in clinical trial outcome prediction. As observed in Table 2, GPT-4 demonstrates a superior performance across most metrics compared to GPT-3.5, suggesting that newer versions of large language models may offer incremental improvements in predicting clinical trial outcomes using the standard prompting method.
| Method | Accuracy | ROC-AUC | PR-AUC | Precision | Recall | F1 |
| GBDT | 0.6250 | 0.8000 | 0.8669 | 0.6250 | 1.0000 | 0.7692 |
| HAtten | 0.7500 | 0.7573 | 0.8718 | 0.8947 | 0.6800 | 0.7727 |
| GPT-3.5 | 0.5250 | 0.4853 | 0.4419 | 0.4000 | 0.5333 | 0.4571 |
| GPT-4 | 0.6500 | 0.6800 | 0.4582 | 0.5385 | 0.4666 | 0.5000 |
| CT-Agent | 0.7000 | 0.8347 | 0.7908 | 0.5714 | 0.8000 | 0.6667 |
4.5 Case Study
@@
We analyze a realistic case study in Table 1. The NCTID is NCT00311402. The trial focused on evaluating the treatment effect of the Aggrenox capsule on cerebrovascular accidents. Aggrenox is a popular drug and contains a combination of aspirin and dipyridamole. First, the user describes the problem using natural language: “I have designed a clinical trial and hope you can help me predict whether this trial can pass” and attach the drug name, disease name, and inclusion/exclusion criteria as features. Then, the Planning Agent decomposes the whole problem into three subproblems based on clinical knowledge. It was enhanced by few-shot learning, which gives some representative examples to the GPT as prompt. These three subproblems are clinical trial enrollment, Aggrenox’s safety to human bodies, and Aggrenox’s efficacy in treating cerebrovascular accidents, which are handled by enrollment agents, safety agents, and efficacy agents, respectively. These agents of subproblems can be solved with the help of external tools. For example, we train an enrollment success prediction model and the estimated enrollment failure probability is 0.359. Enrollment Agent automatically recognizes and call the predictive AI model and insert the results (0.359) into the text. Similarly, Safety Agent (drug safety) identifies that the historical failure rate of Aggrenox capsules is 100%, indicating its high risk. Combining this information, the Reasoning Agent will make the final decision that the trial is highly likely to fail, which is correctly forecasted.
4.6 Ablation Study
4.6.1 Different Versions of GPT
@@
4.6.2 Impact of Few-Shot Learning
This segment of the ablation study investigates the influence of few-shot learning techniques when integrated into CT-Agent, a multi-agent framework employing large language models (LLMs). Our comparative analysis pits the version of the model that incorporates few-shot learning against a baseline version that does not utilize these adaptations. The comparative results are encapsulated in the table below.
| Method | Accuracy | ROC-AUC | PR-AUC | Precision | Recall | F1 |
| CT-Agent w few-shot | 0.7 | 0.8347 | 0.7908 | 0.5714 | 0.8 | 0.6667 |
| CT-Agent w/o few-shot | 0.75 | 0.824 | 0.6793 | 0.647 | 0.7333 | 0.6875 |
Table 3 reveals that while the accuracy and F1 score marginally favor the model without few-shot learning, the few-shot adapted model (CT-Agent) exhibits superior performance in terms of ROC-AUC and PR-AUC. This indicates that the incorporation of few-shot learning significantly bolsters the model’s proficiency in accurately classifying positive instances, despite a minor trade-off in overall accuracy and precision-recall balance.
5 Discussion
This study introduces the groundbreaking Multi-Agent Clinical Trial Helper (CT-Agent), a multi-agent framework that synergizes the advanced capabilities of GPT-4 with sophisticated agent architectures and cutting-edge reasoning technologies like LEAST-TO-MOST and ReAct. Our system significantly enhances the performance of large language models (LLMs) in clinical settings, managing complex trial processes and introducing novel functionalities such as predictive analytics, comprehensive failure analysis, and precise trial duration estimations.
Our evaluations, which include computational benchmarks and expert feedback, underscore the efficiency and effectiveness of CT-Agentin improving clinical trial outcomes. This integration of LLMs with multi-agent systems not only manages the complexities inherent in clinical trials but also bridges the gap between conversational AI and actionable intelligence in healthcare. CT-Agentestablishes a new benchmark in the application of LLMs to clinical trials, promising a future where advanced AI tools play a crucial role in advancing medical research and patient care.
@@ 5 Discussion
While the CT-Agent demonstrates the capability to automatically recognize and decompose user issues, directing them to specialized agents for resolution, it still relies significantly on human intervention for its design and configuration. This dependency on manual input for agent creation limits the system’s scalability and adaptability, particularly in dynamic environments where user requirements and contexts evolve rapidly. Further development could focus on integrating machine learning techniques to enable the CT-Agent to learn from interactions and autonomously update its problem-solving strategies, thereby reducing the need for frequent human oversight and redesign. These points underscore the need for continuous research and development to fully realize the potential of AI-driven clinical trials, addressing both technical and clinical implications.
Appendix A
A.1 Enrollment Model
@@
Embedding Generation: Using the pretrained BioBERT model, the agent converts text data from the criteria, drugs, and diseases into dense vector representations. These embeddings capture deep semantic features that are essential for accurate model predictions.
@@
Evaluation on the test dataset yielded an ROC-AUC score of 0.7037 and an accuracy of 0.7689, indicating effective class differentiation and prediction accuracy. This model exemplifies the integration of advanced NLP and neural architectures to improve clinical trial design and efficiency.
A.2 Function Calling Definitions
@@
"description": "Retrieves information about a drug from the DrugBank database
using the drug’s name as input.",
@@
Given the names of a drug and a disease, the model retrieves
the path connecting the drug to the disease from
the Hetionet Knowledge Graph.
Hetionet is a comprehensive knowledge graph that integrates diverse
biological information by connecting genes, diseases, compounds,
@@
It structures real-world biomedical data into a network,
facilitating advanced analysis and discovery of new insights
into disease mechanisms, drug repurposing,
and the genetic underpinnings of health and disease.",

### v2

@@ 4.2. Experimental Setup
| Method | Accuracy () | ROC-AUC () | PR-AUC () | Precision () | Recall () | F1 () |
| GBDT | 0.6250 | 0.8000 | 0.8669 | 0.6250 | 1.0000 | 0.7692 |
| HAtten | 0.7500 | 0.7573 | 0.8718 | 0.8947 | 0.6800 | 0.7727 |
| GPT-3.5 | 0.5250 | 0.4853 | 0.4419 | 0.4000 | 0.5333 | 0.4571 |
| GPT-4 | 0.6500 | 0.6800 | 0.4582 | 0.5385 | 0.4666 | 0.5000 |
| ClinicalAgent | 0.7000 | 0.8347 | 0.7908 | 0.5714 | 0.8000 | 0.6667 |
| Method | Accuracy () | ROC-AUC () | PR-AUC () | Precision () | Recall () | F1 () |
| ClinicalAgent w few-shot | 0.7 | 0.8347 | 0.7908 | 0.5714 | 0.8 | 0.6667 |
| ClinicalAgent w/o few-shot | 0.75 | 0.824 | 0.6793 | 0.647 | 0.7333 | 0.6875 |
4.3. Implementation Details
@@
Each agent within the ClinicalAgent framework is designated a specific role, which is integrated directly into the LLM’s system prompt for clarity and focus. For instance, the role of the Efficacy Agent is defined as follows:
@@
These implementation strategies collectively ensure that each component of ClinicalAgent operates effectively and that the integration between different agents and external tools is seamless, fostering an environment conducive to robust, reproducible research.
4.4. Quantitative Results
@@
Table 2 presents the performance of various methods. Specifically, ClinicalAgent obtain the highest ROC-AUC score at 0.8347 among all the compared methods. We observe that ClinicalAgent achieve competitive performance among all the well-established methods, e.g., GBDT. Also, compared with the standard prompt (basic GPT) model, our method consistently improves all six evaluation metrics. This result shows that our multi-agent framework brings a significant improvement compared to using GPT directly. We also compare the performance of GPT-3.5 and GPT-4 for the standard prompting method in clinical trial outcome prediction. As observed in Table 2, GPT-4 demonstrates a superior performance across most metrics compared to GPT-3.5, suggesting that newer versions of large language models may offer incremental improvements in predicting clinical trial outcomes using the standard prompting method.
At the same time, we also observed that another critical metric, PR-AUC, while showing significant improvement with ClinicalAgent compared to GPT-Direct method, still lags behind traditional machine learning models. This is particularly important in the healthcare domain, where PR-AUC may hold greater significance. It is noteworthy that in clinical applications, all machine learning models, including HAtten, can be integrated as external tools within the ClinicalAgent framework. In this study, we did not include HAtten as an external tool to focus on comparing the results and highlighting the potential of ClinicalAgent.
4.5. Case Study
@@
We analyze a realistic case study in Table 1 (NCTID: NCT00311402). The trial focused on evaluating the treatment effect of the Aggrenox capsule on cerebrovascular accidents. Aggrenox is a popular drug and contains a combination of aspirin and dipyridamole. First, the user describes the problem using natural language: “I have designed a clinical trial and hope you can help me predict whether this trial can pass” and attach the drug name, disease name, and inclusion/exclusion criteria as features. Then, the Planning Agent decomposes the whole problem into three subproblems based on clinical knowledge. It was enhanced by few-shot learning, which gives some representative examples to the GPT as the prompt. These three subproblems are clinical trial enrollment, Aggrenox’s safety to human bodies, and Aggrenox’s efficacy in treating cerebrovascular accidents, which are handled by enrollment agents, safety agents, and efficacy agents, respectively. These agents of subproblems can be solved with the help of external tools. For example, we train an enrollment success prediction model and the estimated enrollment failure probability is 0.359. Enrollment Agent automatically recognizes and call the predictive AI model and insert the results (0.359) into the text. Similarly, the Safety Agent (drug safety) identifies that the historical failure rate of Aggrenox capsules is 100%, indicating its high risk. Combining this information, the Reasoning Agent will make the final decision that the trial is highly likely to fail, which is correctly forecasted.
4.6. Ablation Study
4.6.1. Different Versions of GPT
@@
4.6.2. Impact of Few-Shot Learning
This segment of the ablation study investigates the influence of few-shot learning techniques when integrated into ClinicalAgent, a multi-agent framework employing large language models (LLMs). Our comparative analysis pits the version of the model that incorporates few-shot learning against a baseline version that does not utilize these adaptations. The comparative results are encapsulated in the table below.
Table 3 reveals that while the accuracy and F1 score marginally favor the model without few-shot learning, the few-shot adapted model (ClinicalAgent) exhibits superior performance in terms of ROC-AUC and PR-AUC. This indicates that the incorporation of few-shot learning significantly bolsters the model’s proficiency in accurately classifying positive instances, despite a minor trade-off in overall accuracy and precision-recall balance.
5. Discussion
This study introduces the groundbreaking Multi-Agent Clinical Trial Helper (ClinicalAgent), a multi-agent framework that synergizes the advanced capabilities of GPT-4 with sophisticated agent architectures and cutting-edge reasoning technologies like LEAST-TO-MOST and ReAct. Our system significantly enhances the performance of large language models (LLMs) in clinical settings, managing complex trial processes and introducing novel functionalities such as predictive analytics, comprehensive failure analysis, and precise trial duration estimations.
Our evaluations, which include computational benchmarks and expert feedback, underscore the efficiency and effectiveness of ClinicalAgent in improving clinical trial outcomes. This integration of LLMs with multi-agent systems not only manages the complexities inherent in clinical trials but also bridges the gap between conversational AI and actionable intelligence in healthcare. ClinicalAgent establishes a new benchmark in the application of LLMs to clinical trials, promising a future where advanced AI tools play a crucial role in advancing medical research and patient care.
@@ 5. Discussion
While the ClinicalAgent demonstrates the capability to automatically recognize and decompose user issues, directing them to specialized agents for resolution, it still relies significantly on human intervention for its design and configuration. This dependency on manual input for agent creation limits the system’s scalability and adaptability, particularly in dynamic environments where user requirements and contexts evolve rapidly. Further development could focus on integrating machine learning techniques to enable the ClinicalAgent to learn from interactions and autonomously update its problem-solving strategies, thereby reducing the need for frequent human oversight and redesign.
These points underscore the need for continuous research and development to fully realize the potential of AI-driven clinical trials, addressing both technical and clinical implications.
Appendix A Supplementary Material
A.1. Enrollment Model
@@
Embedding Generation: Using the pre-trained BioBERT model, the agent converts text data from the criteria, drugs, and diseases into dense vector representations. These embeddings capture deep semantic features that are essential for accurate model predictions.
@@
Evaluation on the test dataset yielded an ROC-AUC score of 0.7037 and an accuracy of 0.7689, indicating effective class differentiation and prediction accuracy. This model exemplifies the integration of advanced Natural Language Processing (NLP) and neural architectures to improve clinical trial design and efficiency.
A.2. Function Calling Definitions
@@
"description": "Retrieves information about a drug
from the DrugBank database using the drug’s
name as input.",
@@
Given the names of a drug and a disease,
the model retrieves the path connecting the drug
to the disease from the Hetionet Knowledge Graph.
Hetionet is a comprehensive knowledge graph that
integrates diverse biological information by
connecting genes, diseases, compounds,
@@
It structures real-world biomedical data into
a network, facilitating advanced analysis and
discovery of new insights into disease mechanisms,
drug repurposing, and the genetic underpinnings
of health and disease.",